### v2

@@ 2.1 Multimodal Multitask Datasets
We target 17 (sub-)tasks related to the LCS process, including 5 tasks for lung nodule detection and characterization, 1 task for cardiovascular disease (CVD) diagnosis, 1 task for CVD mortality prediction, 1 task for lung cancer risk prediction over multiple years, 7 tasks for other chest abnormality exams, 1 task for COVID-19 detection, as well as 1 task for American College of Radiology (ACR) guidelines for Lung CT Screening Reporting and Data System (Lung-RADS) categorization. COVID-19 detection from CT is included since it remains a global threat [55]. The ground-truth labels come from different information sources, including radiology reports, disease history, pathological results, follow-up data, death reports, and laboratory test results for different tasks as described in Figure 3. To curate the multimodal datasets, multiple information sources are aligned, including volumetric CT scans, demographics, smoking history, disease history, cancer history, and family cancer history. In total, 48 different clinical data types are involved in the multimodal datasets, as described in Figure 3. For each task, training, validation, and one or more testing datasets were constructed. An overview of our multimodal multitask dataset is given in Figure 2. The data were collected from different data centers, including NLST, MIDRC, Wake Forest University School of Medicine (WFUSM), and Massachusetts General Hospital (MGH). In total, we curated 17 training, 17 validation, and 33 testing datasets for the 17 tasks, with the detailed composition of each dataset in Table 2. We also collected an out-of-distribution multimodal dataset for transfer learning. To inspect the modeling ability for textural clinical data, we simulated a dataset for clinical information retrieval. Since we unify multitask learning with multimodal data in an MQA framework, each dataset consists of task-specific multimodal inputs, questions, and answers. The details for all tasks are summarised in Table 1.
@@
As the first data source, we were granted access to all recorded data in NLST. NLST is a randomized trial for evaluating lung cancer screening with low-dose CT versus chest radiography. The NLST data were collected from 33 medical institutions, which were randomly indexed without revealing their identifications publicly. The 26,722 participants in the LDCT screening arm were enrolled from August 2002 through April 2004. The participants underwent three screenings at 1-year intervals from August 2002 through September 2007. The follow-up data were collected until December 31, 2009. During the whole process, diverse data were recorded, including demographic characteristics, smoking behavior, disease history, multiple CT series with different reconstruction algorithms and associated imaging parameters, key abnormalities in structured reports, pathological test results for lung cancer, follow-up data, and vital status. Being consistent with clinical practice, we constructed 15 multimodal datasets for 15 tasks, including 5 datasets for predicting the presence of lung nodules and estimating the location, size, margin, and attenuation properties of lung nodules; 7 datasets for identifying chest abnormalities, including atelectasis, segmental or greater, pleural thickening or effusion, non-calcified hilar/mediastinal adenopathy or mass, chest wall abnormality (bone destruction, metastasis, etc.), consolidation, emphysema, reticular/reticulonodular opacities/honeycombing/fibrosis/scar; 1 dataset for cardiovascular abnormality detection; 1 dataset for cardiovascular mortality prediction; and 1 dataset for lung cancer prediction within from 1 to 6 years. Each dataset was randomly split to train, validate, and test datasets. The patient-wise information in the validation and test datasets is not leaked to the training datasets across all tasks. From NLST, we included 125,090 effective volumetric chest CT scans of 26,254 patients.
@@
The second data source is the Medical Imaging and Data Resource Center (MIDRC), a collaboration of leading medical imaging organizations launched in August 2020 as part of NIBIB’s response to the COVID-19 pandemic. We were granted to access all CT series with the associated clinical data. The ground-truth labels for COVID-19 are determined by either the Reverse Transcription Polymerase Chain Reaction (RT-PCR) or rapid antigen test. From MIDRC, we retrieved 35,730 volumetric chest CT series of 7,609 patients being scanned from 2011 to 2021. The dataset was randomly split into the training, validation, and testing datasets.
@@
All CT scans from NLST and MIDRC excluding those in any test datasets were combined as a CT pretraining dataset, comprised of 151,569 CT scans in total. To inspect if the clinical data is effectively encoded, we constructed a clinical question-answering dataset to retrieve key information from the textual clinical data. The integration of all the above-curated datasets is called OpenM3Chest, which will be released upon this paper is accepted. To our best knowledge, OpenM3Chest will be the first-of-its-kind publicly available MQA dataset and also the largest 3D CT dataset for training AI models.
@@
To test the generalizability and adaptability of M3FM, we independently collected two multimodal multitask datasets from WFUSM and MGH, respectively. Each dataset includes CT scans, radiology reports, demographics, smoking history, disease history, personal cancer history, family lung cancer history, and pathology test results for lung cancer. Based on the radiology reports and the pathological test results, we constructed 7 datasets from WFUSM and 6 datasets from MGH for independent evaluation as shown in Figure 2. Specifically, we collected 8,053 patient data at WFUSM from 2015 to 2023, all with radiology reports and 2,000 of them with LDCT and multimodal information. We collected 904 patient data with multimodal information shown in Figure 3 at MGH from 2003 to 2020. The Lung-RADS dataset from WFUSM was randomly split into training, validation, and testing datasets, which were defined to classify the text descriptions into the Lung-RADS category. All other datasets of Wake Forest and MGH were used for testing. To evaluate the transfer learning ability of our M3FM, we collected an out-of-distribution multimodal dataset for non-small cell lung cancer (NSCLC) immunotherapy prognosis from WFUSM. This dataset consists of 90 patient data, including the target label indicating if the patient was diagnosed with immune checkpoint-inhibitor-induced pneumonitis after immunotherapy, the CT scans before immunotherapy, and the clinical variables including the total cycles of Immuno-Oncology (IO), smoking information of pack years, Body Mass Index (BMI) at diagnosis, age, and if the patient received radiation prior to immunotherapy. Among the 90 patients, 49 patients developed immune checkpoint-inhibitor-induced pneumonitis, and the other patients were used as the control group.
@@
Figure 6: Semantic visualization of deep learning-based CVD diagnosis.
Figure 7: Semantic visualization of deep learning-based lung cancer risk prediction.
Table 7: Evaluation results on the MGH datasets in terms of AUC and 95% CI.
Figure 8: Predictions of immunotherapy-induced pneumonitis using different methods.
2.2 M3FM Setting New Records
Figure 4 summarizes the key results of M3FM against the SOTA models [12, 14] on the OpenM3Chest dataset. Overall, M3FM sets new records for all tasks, significantly outperforming the SOTA models for most of these tasks. For a fair comparison, we retrained the Sybil model [12], denoted as $*{}^{*}$, for lung cancer risk prediction without using the costly bounding box annotations but predicting lung cancer risks by merging the separate results of left and right lungs. It is observed that $*{}^{*}$ achieved inferior results for 1$\sim$2-year risk prediction but superior results for 3$\sim$6-year risk prediction in comparison with the results obtained using the original Sybil model. Without using any bounding box, our M3FM achieves AUC of 94.0 (95% 95 CI, 91.19-96.98), 88.81 (95% CI, 85.67-91.95), 85.99 (95% CI, 82.88-89.1), 86.04 (95% CI, 83.1-88.98), 83.92 (95% CI, 80.98-86.85), 82.32 (95% CI, 79.36-85.29) for lung cancer risk prediction over six years, outperforming both $*{}^{*}$ and original Sybil models by the margins of 5% to 9% and 2% to 11%, respectively. For CVD diagnosis and CVD mortality prediction, we compared the results on both the original dataset [14] and our OpenM3Chest dataset. M3FM achieves an AUC of 92.84 (95% CI, 91.36-94.33) for CVD diagnosis and an AUC of 89.04 (95% CI, 84.27-93.81) for CVD mortality prediction on the OpenChest dataset, outperforming the previous models by 5% and 9% respectively, and achieves an AUC of 92.03% for CVD diagnosis and 86.14% for CVD mortality prediction on the datasets constructed in [14], outperforming the previous models by 5% and 10% respectively. For several tasks including nodule detection, nodule localization, nodule size prediction, and emphysema detection, M3FM improves the results by various degrees up to 3% of AUC. For all the other tasks, M3FM significantly improves the performance from $\sim$5to $\sim$10%. To study the scalability of M3FM, we trained three versions of M3FM, consisting of 257M (M3FM-Base), 502M (M3FM-Large), and 865M (M3FM) trainable parameters respectively. The results of these three models are summarized in Figure 5. Overall, with a larger model size, the performance becomes better, especially from M3FM-Base to M3FM-Large. This trend is consistent with the well-known scaling law in the field of foundation models.
2.4 M3FM Identifying Informational Clinical Elements
@@
Then, we evaluated if M3FM models could effectively encode the physical size information. The ablation results in Table 6 show that the embedded physical size information of LDCT improves the AUC results for multiple tasks. The physical size information boosts the AUC of 1$\sim$6-year lung cancer risk prediction by 5%, 4%, 4%, 7%, 8% and 12%, respectively. The physical size information also improves AUC results of the nodule size characterization, CVD diagnosis, and CVD mortality prediction, by 0.71%, 0.47%, and 1.11% respectively.
We quantitatively evaluated the correction of different clinical elements with model outputs by visualizing the attention maps of the last task attention block in M3FM. Figures 6 and 7 respectively visualize the attention heat maps on the selected CT slices and the text tokens of individual patients with CVD or lung cancer risks. In CVD diagnosis, the coronary artery calcification areas were highlighted in the LDCT attention heat maps, and the patients’ disease histories of diabetes, heart disease or heart attack, hypertension, and stroke are highly relevant among text tokens, which is consistent with the quantitative results in Table 4. In predicting lung cancer risks, the lung nodules in LDCT images were localized in the heat maps, and the text tokens related to demographic and family lung cancer histories are more correlated to the model outputs.
@@ 2.5 M3FM Improving Generalizability
We evaluated the generalizability of M3FM models on the independently collected datasets, with the comparison results on the MGH datasets shown in Table 7. For the CVD diagnosis task, we constructed two datasets, which regard moderate and severe CVD as positive and severe CVD only as positive, respectively. On the two CVD datasets, the multimodal multitask model (M3FM-MM-ST) improved the AUC by 10.60% and 6.57% relative to the previous SOTA method, improved the AUC by 4.85% and 2.36% relative to the single-modality single-task model (M3FM-SM-ST), and also achieved slight AUC improvements relative to the multi-modality single-task model M3FM-MM-ST. Relative to M3FM-SM-ST, the M3FM-MM-ST model improved the AUC by 4.39% and 1.75% on the two CVD datasets. For 1-year lung cancer risk prediction, the M3FM-MM-MT model improved the AUC by 20.80% than the previous model under the same experimental settings without using any bounding box annotations, improved the AUC by 4.85% than M3FM-SM-ST, and improved the AUC by 6.89% than M3FM-MM-ST. On the emphysema, atelectasis, and reticular opacities/honeycombing/fibrosis/scar datasets, the M3FM improved the AUC by 14.65%, 14.39%, and 14.80% than the previous model, and also achieved AUC improvements by 0.24% $\sim$ 4.96% than M3FM-SM-ST and M3FM-MM-ST.
2.6 M3FM Enhancing Out-of-Distribution Multimodal Modeling
@@
As a foundational model, we evaluated if M3FM facilitates out-of-distribution multimodal modeling. Here we fine-tuned the M3FM model to predict immunotherapy-induced pneumonitis from volumetric CT prior and the selected clinical data related to immunotherapy as described in Sub-section 2.1. We used the method developed in WFUSM as the baseline and compared different fine-tuned variants of the M3FM model in terms of the average AUC and its standard deviation of five-fold cross-validation. The baseline model used in WFUSM had 89.4% $\pm$ 7.5% AUC by merging all radiomics and clinical features. Specifically, the baseline model was based on a nomogram to predict immunotherapy outcomes using features extracted from radiomic algorithms, a pre-trained ViT-base model, and clinical records. After feature selection, 20 radiomic features, 20 deep features, and 17 clinical features were used for the nomogram. The best result of our fine-tuned M3FM models was 94.1% $\pm$ 2.6% of AUC, which achieved a 4.7% improvement. The M3FM-CT model using CT data only had an AUC of 91.9% $\pm$ 2.6%. The M3FM-Clinical model using clinical text only had a 91.1% $\pm$ 2.9% AUC. The M3FM-Scratch without pretraining achieved 92.5% $\pm$ 2.5% of AUC.

### v3

@@ 2.1 Multimodal Multitask Datasets
Figure 2-(a) shows the general data curation pipeline, including medical tasks of interest definition, task-specific multimodal data collection, multimodal data processing and alignment, and MQA dataset construction. We target 17 (sub-)tasks related to the LCS process, including 5 tasks for lung nodule detection and characterization, 1 task for cardiovascular disease (CVD) diagnosis, 1 task for CVD mortality prediction, 1 task for lung cancer risk prediction over multiple years, 7 tasks for other chest abnormality exams, 1 task for COVID-19 detection, as well as 1 task for American College of Radiology (ACR) guidelines for Lung CT Screening Reporting and Data System (Lung-RADS) categorization. COVID-19 detection from CT is included since it remains a global threat [42]. The ground-truth labels come from different information sources, including radiology reports, disease history, pathology test results, follow-up data, death reports, and laboratory test results for different tasks as described in Table 1. To curate the multimodal datasets, multiple data sources are aligned, including volumetric CT scans, demographics, smoking history, disease history, cancer history, family cancer history, and other task-specific clinical data. In total, 49 different clinical data types are involved in the multimodal datasets for LCS, as described in Table 1. For each task, one training, one validation, and one or more testing datasets were constructed. An overview of our multimodal multitask dataset is given in Figure 2-(b). The data were collected from different data centers and institutes, including NLST, MIDRC, Wake Forest University School of Medicine (WFUSM), and Massachusetts General Hospital (MGH). In total, we curated 17 training, 17 validation, and 29 testing datasets for the 17 tasks, with the detailed composition of each dataset in Figure 2-(c), (d), and (e). We also collected an out-of-distribution multimodal dataset from WFUSM for transfer learning. To inspect the modeling ability for textural clinical data, we simulated a dataset for clinical information retrieval, as illustrated in Figure 1. Since we unify multitask learning with multimodal data in an MQA framework, each dataset consists of task-specific multimodal inputs, questions, and answers. The details for all tasks are summarised in Table 1.
Table 1: Details of the Multimodal Question-Answering Datasets.
Figure 3: (a) Comparison results of the best M3FMs with previous SoTA models in terms of AUC. (b) AUC results with 95% CI of three scales of M3FM models. The AUC results of M3FM and Previous SOTA models are shown. The results of other models can be found in the main text.
@@
As the first data source, we were granted access to all recorded data in NLST NLST is a randomized trial for evaluating LCS with 3D LDCT versus 2D chest radiography, demonstrating that screening with LDCT lowered lung cancer mortality by 20%. The NLST data were collected from 33 medical institutions, which were randomly indexed without revealing their identifications publicly. The 26,722 participants in the LDCT screening arm were enrolled from August 2002 through April 2004. The participants underwent three screenings at 1-year intervals from August 2002 through September 2007. The follow-up data were collected until December 31, 2009. During the whole process, diverse data were recorded, including demographics, smoking history, disease history, multiple CT series with different reconstruction algorithms and associated imaging parameters, key abnormalities in structured reports, pathology test results for lung cancer, follow-up data, and vital status. Being consistent with clinical practice, we constructed 15 multimodal datasets for 15 tasks, including 5 datasets for predicting the presence of lung nodules and estimating the location, size, margin, and attenuation properties of lung nodules; 7 datasets for identifying chest abnormalities, including atelectasis, pleural thickening/effusion, non-calcified hilar/mediastinal adenopathy/mass, chest wall abnormality (bone destruction, metastasis, etc.), consolidation, emphysema, reticular/reticulonodular opacities/honeycombing/fibrosis/scar; 1 dataset for cardiovascular abnormality diagnosis; 1 dataset for cardiovascular mortality risk prediction; and 1 dataset for lung cancer risk prediction within from 1 to 6 years. Each dataset was randomly split into training, validation, and test datasets. The patient-wise information in the validation and test datasets is not leaked to the training datasets across all tasks. From NLST, we included 125,090 effective volumetric chest CT scans of 26,254 patients.
@@
The second data source is the Medical Imaging and Data Resource Center (MIDRC), a collaboration of leading medical imaging organizations launched in August 2020 as part of NIBIB’s response to the COVID-19 pandemic. We were granted to access all CT series with the associated clinical data. The ground-truth labels for COVID-19 are determined by either the Reverse Transcription Polymerase Chain Reaction (RT-PCR) or the Rapid Antigen Test (RAT). From MIDRC, we retrieved 35,730 volumetric chest CT series of 7,609 patients being scanned from 2011 to 2021. The patient data were randomly split into the training, validation, and test datasets.
@@
All CT scans from NLST and MIDRC excluding those in any test datasets were combined as a CT pretraining dataset, comprised of 151,569 CT scans in total. To inspect if the clinical data are effectively encoded, we constructed a clinical question-answering dataset to retrieve key information from the textual clinical data. The integration of all the above-curated datasets is called OpenM3Chest, which will be released upon this paper is accepted. To our best knowledge, OpenM3Chest will be the first-of-its-kind publicly available MQA dataset and also the largest 3D CT dataset for training AI models.
@@
To test the generalizability and adaptability of M3FM, we independently collected two multimodal multitask datasets from the third and fourth data sources, i.e., WFUSM and MGH, respectively. Each dataset includes CT scans, radiology reports, demographics, smoking history, disease history, personal cancer history, family lung cancer history, and pathology test results for lung cancer. The MGH and WFUSM review boards approved the analysis of all these multimodal data and tasks. Based on the radiology reports and the pathology test results, we constructed 7 datasets from WFUSM and 6 datasets from MGH for independent evaluation as shown in Figure 2. Specifically, we collected 8,053 patient data at WFUSM from 2015 to 2023, all with radiology reports and 1,800 of them with LDCT and multimodal information. We collected 904 patient data with multimodal data shown in Table 1 at MGH from 2016 to 2021. The Lung-RADS dataset from WFUSM was randomly split into training, validation, and test datasets, which were defined to classify the text descriptions into the Lung-RADS category. All other datasets of WFUSM and MGH were used for testing. To evaluate the adaptability of our M3FM, we collected an out-of-distribution multimodal dataset for non-small cell lung cancer (NSCLC) immunotherapy prognosis from WFUSM. This dataset consists of 90 patient data, including the target label indicating if the patient was diagnosed with immune checkpoint-inhibitor-induced pneumonitis after immunotherapy, the CT scans before immunotherapy, and the clinical variables including the total cycles of Immuno-Oncology (IO), smoking information of pack years, Body Mass Index (BMI) at diagnosis, age, and if the patient received radiation prior to immunotherapy. Among the 90 patients, 49 patients developed immune checkpoint-inhibitor-induced pneumonitis, and the other patients were used as the control group.
The details on the multimodal data processing and alignment and the MQA dataset construction are described in Section 4.1.
2.2 M3FM Setting New Records
Figure 3-(a) summarizes the key results of M3FM against the SOTA models [33, 35] on the OpenM3Chest dataset. Overall, M3FM sets new records for all tasks, significantly outperforming the SOTA models for most of these tasks. For a fair comparison, we retrained the Sybil model [33], denoted as Sybil, for lung cancer risk prediction without using the costly bounding box annotations but predicting lung cancer risks by merging the separate results of left and right lungs. It is observed that Sybil achieved inferior results for 1$\sim$2-year risk prediction but superior results for 3$\sim$6-year risk prediction in comparison with the results obtained using the original Sybil model. Without using any bounding box, our M3FM achieves an Area Under the receiver operating characteristic Curve (AUC) of 94.0 (95% 95 CI, 91.19-96.98), 88.81 (95% CI, 85.67-91.95), 85.99 (95% CI, 82.88-89.1), 86.04 (95% CI, 83.1-88.98), 83.92 (95% CI, 80.98-86.85), 82.32 (95% CI, 79.36-85.29) for lung cancer risk prediction over six years, outperforming both Sybil and original Sybil models by the margins of 5% to 9% and 2% to 11%, respectively. For CVD diagnosis and CVD mortality prediction, we compared the results on both the original dataset [35] and our OpenM3Chest dataset. M3FM achieves an AUC of 92.84 (95% CI, 91.36-94.33) for CVD diagnosis and an AUC of 89.04 (95% CI, 84.27-93.81) for CVD mortality prediction on the OpenChest dataset, outperforming the previous models (Tri2D-Net [35]) by 5% and 9% respectively, and achieves an AUC of 92.03% for CVD diagnosis and 86.14% for CVD mortality prediction on the datasets constructed in [35], outperforming the previous models (Tri2D-Net) by 5% and 10% respectively. For several tasks including nodule detection, nodule localization, nodule size prediction, and emphysema detection, M3FM improves the results by various degrees up to 3% of AUC. For all the other tasks, M3FM significantly improves the performance from $\sim$5to $\sim$10%. To study the scalability of M3FM, we trained three versions of M3FM, consisting of 257M (M3FM-Base), 502M (M3FM-Large), and 865M (M3FM-Huge) trainable parameters respectively. The results of these three models are summarized in Figure 3-(b). Overall, with a larger model size, the performance becomes better, especially from M3FM-Base to M3FM-Large. This trend is consistent with the well-known scaling law in the field of foundation models.
@@ 2.3 M3FM Encoding Multimodal Data and Synergizing Multiple Clinical Tasks
Figure 4: Qualitative inspection of the task encoder. (a) Attention map visualization of the task encoder for two CVD diagnosis examples. (b) Attention map visualization of the task encoder for two lung cancer risk prediction examples.
2.4 M3FM Identifying Clinically Informational Elements
@@
Then, we evaluated if M3FMs could effectively encode the physical size information. The ablation results in Table 5 show that the embedded physical size information of LDCT improves the AUC results for multiple tasks. The physical size information boosts the AUC of 1$\sim$6-year lung cancer risk prediction by 5%, 4%, 4%, 7%, 8% and 12%, respectively. The physical size information also improves AUC results of the nodule size characterization, CVD diagnosis, and CVD mortality prediction, by 0.71%, 0.47%, and 1.11% respectively.
We quantitatively evaluated the correction of different clinical elements with model outputs by visualizing the attention maps of the last task attention block in M3FM. Figure 4 visualizes the attention heat maps on the selected CT slices and the text tokens of individual patients with CVD or lung cancer risks. In CVD diagnosis, the coronary artery calcification areas were highlighted in the LDCT attention heat maps, and the patients’ disease histories of diabetes, heart disease or heart attack, hypertension, and stroke are highly relevant among text tokens, which is consistent with the quantitative results in Table 3. In predicting lung cancer risks, the lung nodules in LDCT images were localized in the heat maps, and the text tokens related to demographic and family lung cancer histories are more correlated to the model outputs.
Table 6: Evaluation results on the MGH and WFUSM datasets in terms of AUC and 95% CI.
@@ 2.5 M3FM Improving Generalizability
We evaluated the generalizability of M3FMs on the independently collected datasets, with the comparison results on the MGH and WFUSM datasets shown in Table 6. For the CVD diagnosis task, we constructed two datasets, which regard moderate and severe CVD as positive and severe CVD only as positive, respectively. On the two MGH CVD datasets, the multimodal multitask model (M3FM-MM-MT) improved the AUC by 10.60% and 6.57% relative to the previous model, improved the AUC by 4.85% and 2.36% relative to the single-modality single-task model (M3FM-SM-ST), and also achieved slight AUC improvements relative to the multi-modality single-task model (M3FM-MM-ST). Relative to M3FM-SM-ST, the M3FM-MM-ST model improved the AUC by 4.39% and 1.75% on the two CVD MGH datasets. For the 1-year lung cancer risk prediction on the MGH dataset, the M3FM-MM-MT model improved the AUC by 20.80% than the previous model under the same experimental settings without using any bounding box annotations, improved the AUC by 4.85% than M3FM-SM-ST, and improved the AUC by 6.89% than M3FM-MM-ST. On the MGH emphysema, atelectasis, and reticular opacities/honeycombing/fibrosis/scar datasets, the M3FM improved the AUC by 5.23%, 14.34%, and 12.91% than the previous model, and also achieved AUC improvements by 0.24% $\sim$ 4.96% than M3FM-SM-ST and M3FM-MM-ST. For the CVD tasks on MGH datasets, M3FM-MM-MT improved the AUC by 12% and 6.29% than the previous model, improved the AUC by 6.46% and 3.77% relative to M3FM-SM-ST, and had the same results with M3FM-MM-ST. For the 1-year lung cancer risk prediction on the MGH dataset, the M3FM-MM-MT model improved the AUC by 18.54% relative to the previous model under the same experimental settings without using any bounding box annotations, improved the AUC by 5.91% than M3FM-SM-ST, and improved the AUC by 2.57% than M3FM-MM-ST; and M3FM-MM-ST improved the AUC by 3.24% than M3FM-SM-ST. On the WFUSM emphysema, atelectasis, and reticular opacities/honeycombing/fibrosis/scar datasets, the M3FM-MM-MT model improved the AUC by 0.78%, 9.65%, and 14.24% than the previous model.
Figure 5: (a) The same M3FM architecture was fine-tuned to perform the new immunotherapy prognosis task with different CT and clinical inputs. (b) Results of immunotherapy-induced pneumonitis using different methods.
2.6 M3FM Enhancing Out-of-Distribution Multimodal Analysis
@@
As a foundational model, we evaluated if M3FM facilitates out-of-distribution multimodal modeling. Here we fine-tuned the M3FM to predict immunotherapy-induced pneumonitis from volumetric CT prior and the selected clinical data related to immunotherapy as described in Sub-section 2.1. We used the method developed in WFUSM as the baseline and compared different fine-tuned variants of the M3FM in terms of the average AUC and its standard deviation of five-fold cross-validation. The baseline model used in WFUSM had 89.4% $\pm$ 7.5% AUC by merging all radiomics and clinical features. Specifically, the baseline model was based on a nomogram to predict immunotherapy outcomes using features extracted from radiomic algorithms, a pre-trained ViT-base model, and clinical records. After feature selection, 20 radiomic features, 20 deep features, and 17 clinical features were used for the nomogram. The best result of our fine-tuned M3FMs was 94.1% $\pm$ 2.6% of AUC, which achieved a 4.7% improvement. The M3FM-CT model using CT data only had an AUC of 91.9% $\pm$ 2.6%. The M3FM-Clinical model using clinical text only had a 91.1% $\pm$ 2.9% AUC. The M3FM-Scratch without pretraining achieved 92.5% $\pm$ 2.5% of AUC.